Hydrates:

1a. Calcium chloride (CaCl2) is available; the procedure calls for an 8% CaCl2•10 H20 solution. What would be the procedure for making 2,000 mL of the desired concentration?

1b. How many grams of CuSO4•5 H20 are needed to make 1 L of a 1% solution of CuSO4?

1c. How many grams of Co(NO3)2•6 H2O are needed in preparing 400 mL of a 0.02% solution of Co(NO3)2?

1d. A 9% CuSO4 (anhydrous) solution is made up by dissolving 9 g of anhydrous CuSO4 in 100 mL of solution. How is a 9% CuSO4•5 H2O solution made?
Answer #1

1 (a) Mass of CaCl2. 8 H2O to be dissolved = 8 * 2000 /100 grams = 160 grams

You have to weigh 160 grams of CaCl2. 8 H2O and dissolve it in water to make a solution of 2000 mL.

1 (b) Mass of CuSO4. 5 H2O = 1* 1000/ 100 grams = 10 grams

1 (c) Mass of Co(NO3)2 = 0.02 * 400 / 100 grams = 0.08 grams

1 (d) By dissolving 9 grams of CuSO4.5H2O in water to make 100 mL of the solution.